find the slope-intercept equation of the line with the following properties perpendicular to the line x-7y=5; containing the point (3,3).
Answer by TimothyLamb(4379)

---
given:
(3,3)
x-7y=5
---
7y = x - 5
y = (1/7)x - 5/7
---
perp slope = -7
---
to solve the linear equation, copy and paste this (the given point and perpendicular slope):
3, 3, -7
into the "Point-Slope form: x1 y1 m" input box here: https://sooeet.com/math/linear-equation-solver.php
---
x-intercept = (3.42857143, 0)
y-intercept = (0, 24)
slope = -7
---
answer:
the slope-intercept form of our new line:
y = -7x + 24
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php